OpenStudy (anonymous):

Which of the following is a sinusoid? y=sinx x^2+y^2=1 y=^3sqrtx y=[x]

OpenStudy (eric_d):

A sine wave, or sinusoid, is the graph of the sine function in trigonometry. A sinusoid is the name given to any curve that can be written in the form (A and B are positive). Sinusoids are considered to be the general form of the sine function.

OpenStudy (anonymous):

so which one??

OpenStudy (eric_d):

It should be y=sinx

OpenStudy (eric_d):

The graphs of y = sin(x) and y = cos (x) are called as sinusoidal graphs. In both these graphs, the shape of the graph begins repeating after an interval of 2π. In fact, any co terminal angles will have the same sin(x) and cos (x) values. That is, if the graphs of y = sin(x) and y = cos (x) is shifted horizontally by 2π, the resulting shape would be identical to the original graph. Sinusoidal functions are a special case of periodic function!